OpenStudy (anonymous):

Happy Feet Rentals charges $14 per day for snorkel gear and $16 per day to rent a wetsuit. Sam and Monica paid a total of $300 to rent a set of snorkel gear and a wetsuit each for a few days. Monica rented them for 2 days less than Sam. The following equation represents this situation, where x represents the number of days Sam rented a surfboard and a wetsuit. 14x + 16x + 14(x - 2) + 16(x - 2) = 300 What is the total amount Sam paid to rent the snorkel gear and wetsuit?

OpenStudy (anonymous):

you dont have to write it out... im on a time limit and its cuttin deep now... this determines if i pass my grade if you have the answer, i will give out all medals . please

OpenStudy (anonymous):

14x + 16x + 14(x - 2) + 16(x - 2) = 300 30x + 30x -28 - 32 = 300 60x = 300 - 623 =>x = 3.97

OpenStudy (anonymous):

Oh, why don't you say so :)

OpenStudy (anonymous):

$240 $180 $80 $150

OpenStudy (anonymous):

those are the options

OpenStudy (anonymous):

180

OpenStudy (anonymous):

4x + 16x + 14(x - 2) + 16(x - 2) = 300 30x + 30x -28 - 32 = 300 60x = 300 +60 x=6 therefore surfboard: = 6*14 =84 wetsuit = 6*16 = 96 Total paid by Sam = 96 + 84 = 180

OpenStudy (anonymous):

Chil, you error was on line 3. -28-32 = -60 so add 60 on both sides therfore 60x = 300 +60 = 360 => x = 6.

OpenStudy (anonymous):

@chemwile Thanks, I'm sort of running back forth many places!